OpenStudy (anonymous):

What is the simplified form of (8x/x^2-4)/(4x^2/x^2+4x+4)

jhonyy9 (jhonyy9):

us this a/b/c/d = a/b * d/c ok ?

OpenStudy (anonymous):

what ?

jhonyy9 (jhonyy9):

so how you see there when you divide two fractions so than you need multiplie first by second inversed ok ?

jhonyy9 (jhonyy9):

so than how will be this your exercise ?

jhonyy9 (jhonyy9):

you like make the calcule ?

jhonyy9 (jhonyy9):

so than will be 8x x^2 +4x +4 ------- * ------------ = x^2 -4 4x^2 can you simplifie it now ?

OpenStudy (anonymous):

i have it